Supposed the random variable X has a normal distribution with a mean of 10 and variance of 10 .Calculate p(z>20)

Continues  probability 
Answer #1

Solution :

Given that ,

mean = \mu = 20

variance = 9

standard deviation = \sigma = \sqrt 9 = 3

X\sim N(\mu = 20 , \sigma = 3)

a)

P(X \leq 10) = P((X - \mu ) / \sigma \leq (10 - 20) / 3)

= P(z \leq -3.33)   

Using standard normal table,

=0.0004   

Probability = 0.0004

b)

P(15 \leq X \leq 20) = P((15 - 20)/ 3 \leq (X - \mu ) / \sigma \leq (20 - 20) / 3 )

= P( -1.67 \leq z \leq 0)

= P(z \leq 0) - P(z \leq -1.67)

Using standard normal table,  

0.5 - 0.0475

= 0.4525

Probability = 0.4525

c)

Given that ,

mean = \mu = 2

variance = 12

standard deviation = \sigma =\sqrt12 = 3.4641

Y\sim N(\mu = 2 , \sigma = 3.4641)

Z = 2X + Y + 3

E(Z) = E( 2X + Y + 3)

= E(2X) + E(Y) + 3 by additive property of expectation

=2E(X) + E(Y) + 3

= 2*20 + 2 + 3

= 40+5

= 45

E(Z) = 45

V(Z) = V(2X + Y + 3)

= V(2X) + V(Y) +0 , by additive property of variance.

Since X and Y are independent to each other.

= 4V(X) + 12

=4*9 + 12

=36+12

= 48

V(Z) = 48


answered by: ANURANJAN SARSAM